the speed of a boat in still water is 24 miles per hour. if the boat travels 54 miles upstream in the same time it takes to travel 90 miles downstream, what is the rate of the stream?
Found 2 solutions by Paul, checkley71:
Answer by Paul(988)

Let the speed of the current be x
SPeed with the water -->24+x
Speed aginst the water --->24-x
Equation:

Cross multiply 
54(24+x)=90(24-x)
1296+54x=2160-90x
144x=864
x=6

Hence, the speed of the current is 6mph.
Paul.

Answer by checkley71(8403)

54/(24-C)=90/(24+C) CROSS MULTIPLYING GIVES 1296+54C=2160-90C OR 144C=864 OR
C=6 THEREFORE THE CURRENT IS 6MPH
PROOF 54/(24-6)=90/(24+6) OR 54/18=90/30 PR 3=3